The Magazine Mass Marketing Company has received 15 entries in its latest sweepstakes. They know that the probability of receiving a magazine subscription order with an entry form is 0.6 . What is the probability that no less than 4 of the entry forms will include an order? Round your answer to four decimal places.
Answer #1

Binomial distribution: P(X) = nCx px qn-x

Number of entries, n = 15

P(order), p = 0.6

q = 1 - p = 0.4

P(no less than 4 of the entry forms will include an order) = P(X 4)

= 1 - P(X < 4)

= 1 - [P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3)]

= 1 - (0.415 + 15x0.6x0.414 + 15C2x0.62x0.613 + 15C3x0.63x0.412)

= 1 - 0.0019

= 0.9981